OpenStudy (anonymous):

Solve: 5(x + 6) Less than or equal to 3x - 4 x => -17 x =>17 x <= -17 x <=-5

OpenStudy (ryan123345):

\[5(x+6)\le3x-4\]

OpenStudy (anonymous):

@skullpatrol -_____- Stop commenting on my questions.

OpenStudy (ryan123345):

\[5x+30\le3x-4\]

OpenStudy (ryan123345):

\[2x+30\le-4\]

OpenStudy (anonymous):

@ryan123345 Thank you for actually helping me :)

OpenStudy (ryan123345):

\[2x \le-34\]

OpenStudy (ryan123345):

\[x \le-17\]

OpenStudy (ryan123345):

Your welcome :) and @skullpatrol leave her alone.

OpenStudy (ryan123345):

dude just stop

OpenStudy (anonymous):

Thank you :D

OpenStudy (ryan123345):

your welcome. and skull patrol im giving u one more chance. If u comment one more time ur done :)

OpenStudy (ryan123345):

@karatechopper @AccessDenied

OpenStudy (ryan123345):

Can you please tell this guy to stop bothering us? @karatechopper @AccessDenied thank u :)